OpenStudy (anonymous):

the fibbonacci sequence is defined by an+1 = an+1 +an for n more than and equal to 1 where a1= 1, a2=2 find limit an+1/an assuming it exists

OpenStudy (anonymous):

oh i see, it is \[\lim_{n\to \infty}\frac{a_{n+1}}{a_n}\] a famous number

OpenStudy (anonymous):

yup but i did not understand this chapter much in the class..can you help me?

OpenStudy (anonymous):

sure i think

OpenStudy (anonymous):

rewrite the numerator \(a_n+a_{n-1}\)

OpenStudy (anonymous):

\[\lim_{n\to \infty}\frac{a_{n+1}}{a_n}=\lim_{n\to \infty}\frac{a_n+a_{n-1}}{a_n}\]

OpenStudy (anonymous):

first off is that clear?

OpenStudy (anonymous):

sorry..i forgot to give the hint..an+1/an = an+2/an+1

OpenStudy (anonymous):

no need

OpenStudy (anonymous):

is it clear what i wrote?

OpenStudy (anonymous):

that is just step one we do a small bit of algebra, then solve a quadratic equation it will not take long

OpenStudy (anonymous):

where did u get an + an - 1?

OpenStudy (anonymous):

that is how the fibonacci sequence is defined each term is the sum of the two preceding terms so for example \[a_8=a_7+a_6\] and \[a_{101}=a_{100}+a_{99}\] and \[a_{n+1}=a_n+a_{n-1}\]

OpenStudy (anonymous):

clear or no?

OpenStudy (anonymous):

clear :)

OpenStudy (anonymous):

ok so now we are at this step \[\lim_{n\to \infty}\frac{a_n+a_{n-1}}{a_n}\] we we break in to two parts \[\lim_{n\to \infty}1+\frac{a_{n-1}}{a_n}\]

OpenStudy (anonymous):

let me now if that is ok, it is algebra

OpenStudy (anonymous):

ok i got it..next?

OpenStudy (anonymous):

ok now comes a small thinking part

OpenStudy (anonymous):

\[\frac{a_{n+1}}{a_n}\] is the next term over the previous one, wheras \[\frac{a_{n-1}}{a_n}\] is the previous term over the next term

OpenStudy (anonymous):

clear?

OpenStudy (anonymous):

ok clear

OpenStudy (anonymous):

so if \[\lim_{n\to \infty}\frac{a_{n+1}}{a_n}=x\] (here we assume it exists) then \[\lim_{n\to \infty}\frac{a_{n-1}}{a_n}=\frac{1}{x}\]

OpenStudy (anonymous):

as one is the reciprocal of the other, at least in the limit

OpenStudy (anonymous):

only one step to go so far so good?

OpenStudy (anonymous):

no..i don't understand at 1/ x

OpenStudy (anonymous):

ok that is really the only part that requires thought, so i have to say it in english

OpenStudy (anonymous):

\[\frac{a_{n}}{a_{n+1}}\] is the reciprocal of \[\frac{a_{n+1}}{a_n}\] should be clear

OpenStudy (anonymous):

now as you take the limit as n goes to infinity there is no difference between \[\lim_{n\to \infty}\frac{a_n}{a_{n+1}}\] and \[\lim_{n\to \infty}\frac{a_{n-1}}{a_n}\]

OpenStudy (anonymous):

what is says in english is the limit of the previous term over the next term

OpenStudy (anonymous):

the subscripts are unimportant except that the indicate that you are taking the limit of the ration of one term over the previous one i hope that is clear

OpenStudy (anonymous):

ohhh i see..there is no difference as we use infinity right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

you replace \(n\) by \(n+1\) everywhere and nothing changes

OpenStudy (anonymous):

ok so now what do we have? putting it together we have \[\lim_{n\to \infty}\frac{a_{n+1}}{a_n}=1+\lim_{n\to \infty}\frac{a_{n-1}}{a_n}\] and assuming the limit exists and is say \(x\) this equation becomes \[x=1+\frac{1}{x}\]

OpenStudy (anonymous):

then do quadratic right?

OpenStudy (anonymous):

and now i have to run, but this is a simple enough equation to solve you get \[x^2=x+1\] or \[x^2-x-1=0\] which will solve using the quadratic formula, get a famous number called "the golden ration" sometimes denoted as \(\phi\)

OpenStudy (anonymous):

yeah, a quadratic actually you get two answers, pick the one that is greater than one, since the limit obviously is google "golden ratio" and you will get more information than you can use

OpenStudy (anonymous):

ok i got it..thank you very very much for helping..

OpenStudy (anonymous):

yw